OpenStudy (anonymous):

Y/5 + 10 = -15 whats y?

OpenStudy (anonymous):

Well, first you need to get Y by itself, so subtract 10. What is -15 010? -25 right?

OpenStudy (anonymous):

Sorry that was supposed to say -15 -10 CX So then you have y/5=-25 Now multiply both sides by the reciprocal of 1/5, which is 5. So you have y=-25 x 5 Which would be -125 :) Now pug that in and see if it makes sense.

OpenStudy (anonymous):

-125/5=-15

OpenStudy (unanimoose):

y/5 + 10 = -15 Subtract 10 from both sides. y/5 + 10 - 10 = -15 - 10 y/5 = -25 Now multiply both sides by 5. y/5 * 5 = -25 * 5 y = -125 Now check. -125/5 + 10 = -15 -25 + 10 = -15 -15 = -15 What Kit Kat said. LOL

OpenStudy (anonymous):

LOL CX Good job, @Unanimoose CX

OpenStudy (unanimoose):

I didn't want it to go to waste so I posted it anyway. xD

OpenStudy (anonymous):

xD Well, now she has two sources verifying that the answer is =125 :)

OpenStudy (unanimoose):

LOL yeah. :P